OpenStudy (anonymous):

I need help solving 1/x+1/x+4=1/4

OpenStudy (anonymous):

Welcome to Openstudy!

OpenStudy (anonymous):

\[\Large \frac{ 1 }{ x }+\frac{ 1 }{ x+4 }=\frac{ 1 }{ 4 }\] Is this your question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the lcd is 4x, right? but after that I am lost

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

No it's not. With a question like this, it is very tedious yes, but the lcd (when you have two unlike denominators with variables) is the product of the two

OpenStudy (anonymous):

so lcd=x(x+4)

OpenStudy (anonymous):

\[\Large [x(x+4)] \times\frac{ 1 }{ x }+\frac{ 1 }{ x+4 }=\frac{ 1 }{ 4 } \times [x(x+4)]\]

OpenStudy (anonymous):

You could add a 4 in there to avoid one step, doesn't really matter

OpenStudy (anonymous):

\[\Large (x+4)+(x)=\frac{[x(x+4)]}{4}\]

OpenStudy (anonymous):

multiplying both sides by 4

OpenStudy (anonymous):

\[\Large 4[(x+4)+x]=x(x+4)\]

OpenStudy (anonymous):

ok I factored to get: 8x+16= x^2 +4x

OpenStudy (anonymous):

That's correct

OpenStudy (anonymous):

Go on from there

OpenStudy (anonymous):

x^2-4x-16=0 is what you end up with

OpenStudy (anonymous):

will i make this equation equal zero?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you gotta use the quadratic equation: http://www.math.com/students/calculators/source/quadratic.htm

OpenStudy (anonymous):

where a=1 b=-4 and c=-16

OpenStudy (anonymous):

i figured that bc I couldn't factor it out. okay

OpenStudy (anonymous):

[2-\sqrt{20}, 2+\sqrt{20}\]

OpenStudy (anonymous):